### v1

@@ 1.2 Contributions
We propose a Lipschitz-free cubic regularization (LF-CR) algorithm for solving convex-concave minimax optimization problems without the knowledge of the Lipschitz constants. It can be proved that the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution with respect to the restricted primal-dual gap is upper bounded by $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{3}}{\epsilon})^{\frac{2}{3}}$, where $z^{0}=(x^{0},y^{0})$ is a pair of initial points, $z^{*}=(x^{*},y^{*})$ is a pair of the optimal solutions, and $\rho$ is the Lipschitz constant. This iteration complexity has been proved to be optimal for solving convex-concave minimax optimization problems [46, 12]. Although the adaptive second-order optimistic method in [40] has the same optimal complexity in terms of the restricted primal-dual gap function, an additional assumption of Lipschitz continuous gradient is needed.
@@
We further propose a fully parameter-free cubic regularization (FF-CR) algorithm without any knowledge of the problem parameters, including the Lipschitz constants and the upper bound of the distance from the initial point to an optimal solution. The iteration complexity to obtain an $\epsilon$-optimal solution with respect to the gradient norm can be proved to be upper bounded by $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{2}}{\epsilon})^{\frac{2}{3}}$. Numerical experiments show the efficiency of the two proposed algorithms.
@@
To the best of our knowledge, the proposed FF-CR algorithm is the first fully parameter-free second-order algorithm for convex-concave minimax optimization problems. Moreover, the complexity of the algorithm is consistent with the optimal iteration complexity lower bound of the existing second-order algorithm under the assumption of knowing the parameters for solving convex-concave minimax problems [46, 12]. Compared to the state-of-the-art algorithm with the knowledge of parameters in [34], we improve the complexity for the convex-concave setting by a logarithmic factor. Moreover, our method achieves better iteration complexity compared to the adaptive second-order optimistic parameter free method in [40], which has an iteration complexity of $\mathcal{O}(\epsilon^{-1})$ in terms of gradient norm. Furthermore, our method requires only a mild assumption that the Hessian is Lipschitz continuous, without making additional assumptions as in [25, 34, 40].
@@ 1.3 Organization
$DF(z)=\begin{bmatrix}~{}~{}\nabla_{xx}^{2}f(x,y)&~{}~{}\nabla_{xy}^{2}f(x,y)\\ -\nabla_{yx}^{2}f(x,y)&-\nabla_{yy}^{2}f(x,y)\end{bmatrix}\in\mathbb{R}^{(m+n)\times(m+n)}.$ (3)
@@ 2 A Lipschitz-Free Cubic Regularization (LF-CR) Algorithm
In this section, we propose a Lipschitz-free cubic regularization (LF-CR) algorithm that does not require prior knowledge of the Lipschitz constant. We iteratively approximate the Lipschitz constants and show that the iteration complexity of the proposed algorithm is still optimal in the absence of information about the Lipschitz constants. Moreover, the LF-CR algorithm is a key subroutine in our fully parameter-free algorithm that we will present in the next section.
The proposed LF-CR algorithm is based on the Newton-MinMax algorithm [47], which consists of the following two important algorithmic components at each iteration:
@@
    $\hat{z}_{k+1}=\hat{z}_{k}-\lambda_{k+1}F(z_{k+1}).$
Although the Newton-MinMax algorithm achieves the optimal iteration complexity, it requires knowing $\rho$, the Lipschitz constant of the Hessian of $f(x,y)$, which limits its applicability. To eliminate the need for the prior knowledge of $\rho$, we propose a LF-CR algorithm that employs a line search strategy in the first gradient update step. Similar line search strategies have been used in [10, 52, 55, 56]. More detailedly, in each iteration of the proposed LF-CR algorithm, we aim to search a pair $(H_{k,i},z_{k,i})$ by backtracking such that
where $z_{k,i}$ is a solution of the following nonlinear equation:
We then introduce how to solve the nonlinear equations in (6). Similar to the discussion in [47], (6) can be reformulated as finding a pair of $(\theta_{k,i},z_{k,i})$ such that
$\theta_{k,i}$ $=6H_{k,i}\|z_{k,i}-\hat{z}_{k}\|,$
$(DF(\hat{z}_{k})$ $+\theta_{k,i}I_{m+n})(z_{k,i}-\hat{z}_{k})=-F(\hat{z}_{k}).$
Solving the above nonlinear system of equations is equivalent to first solving the following univariate nonlinear equation about $\theta$:
$\phi(\theta):=\sqrt{\|(DF(\hat{z}_{k})+\theta I_{m+n})^{-1}F(\hat{z}_{k})\|^{2}}-\frac{\theta}{6H_{k,i}}=0.$ (7)
Similar to the proof of Proposition 4.5 in [47], when $\theta>0$, $\phi(\theta)$ is a strictly decreasing and convex function. Therefore, we can first use Newton’s method to solve the one-dimensional nonlinear equation (7) to find the optimal solution of $\theta$. Then by substituting it into the following equation and solving the following linear system,
@@
we obtain $z_{k,i}$. Next, the proposed LF-CR algorithm performs an extra gradient update on $z_{k,i}$. The detailed algorithm is formally described in Algorithm 1.
In the following subsection, we establish the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the restricted primal-dual gap.
@@
Step 1: Input $x_{0},y_{0},H_{0}$; Set $k=1$, $\hat{x}_{1}=x_{1}=x_{0},\hat{y}_{1}=y_{1}=y_{0}$.
@@
   $H_{k}=H_{k,i}$, $z_{k+1}=z_{k,i}$; otherwise, set $H_{k,i+1}=2H_{k,i}$, $i=i+1,$ go to Step
   2(b).
Step 3: Update $\hat{z}_{k}$:
$\hat{z}_{k+1}=\hat{z}_{k}-\lambda_{k+1}F(z_{k+1}).$ (11)
Step 4: If some stationary condition or stopping criterion is satisfied, let $\bar{z}_{k}=\frac{1}{\sum_{i=1}^{k}\omega_{i+1}}\sum_{i=1}^{k}\omega_{i+1}z_{i+1}$ and output $(\bar{z}_{k},H_{k})$; otherwise, set $k=k+1,$ go to Step 2(a).
@@ 2.1 Complexity analysis
Before we prove the iteration complexity of the LF-CR algorithm, we first define the following stationarity gap function, which provides a performance measure for how close a point $(\hat{x},\hat{y})$ is to $(x^{*},y^{*})$, which is an optimal solution of (1) under convex-concave setting. Similar to [47], we first define the restricted primal-dual gap function for (1) as follows.
Definition 1.
@@
$\textsc{gap}(\hat{x},\hat{y};\beta)=\max_{y\in\mathbb{B}_{\beta}(y^{\star})}f(\hat{x},y)-\min_{x\in\mathbb{B}_{\beta}(x^{\star})}f(x,\hat{y}),$ (12)
where $(x^{*},y^{*})$ is a saddle point of $f(x,y)$ and $\beta$ is large enough so that for a given $(\hat{x},\hat{y})$, $\|\hat{z}-z^{\star}\|\leqslant\beta$, $\mathbb{B}_{\beta}(x^{*})=\{x:\|x-x^{*}\|\leqslant\beta\}$ and $\mathbb{B}_{\beta}(y^{*})=\{y:\|y-y^{*}\|\leqslant\beta\}$.
Definition 2.
A pair $(\hat{x},\hat{y})$ is called an $\epsilon$-optimal solution of (1) with respect to the restricted primal-dual gap if $\textsc{gap}(\hat{x},\hat{y};\beta)\leqslant\epsilon.$
We also need to make the following assumption about the smoothness of $f(x,y).$
Assumption 1.
$f(x,y)$ is Hessian Lipschitz continuous, that is, there exists a positive scalar $\rho$ such that for any $(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$,
$\|\nabla^{2}f(x,y)-\nabla^{2}f(x^{\prime},y^{\prime})\|\leqslant\rho~{}\|(x-x^{\prime},y-y^{\prime})\|.$
@@
Lemma 1.
Suppose that Assumption 1 holds. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be sequences generated by Algorithm 1. Set $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then for any $z\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ and $s\geqslant 1$, we have
@@
Proof.
By (11) in Algorithm 1, we have
$~{}\lambda_{k+1}\langle z_{k+1}-z,F(z_{k+1})\rangle$
$=$ $~{}\lambda_{k+1}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle+\lambda_{k+1}\langle\hat{z}_{k+1}-z_{0},F(z_{k+1})\rangle+\lambda_{k+1}\langle z_{0}-z,F(z_{k+1})\rangle$
$=$ $~{}\lambda_{k+1}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle+\langle\hat{z}_{k+1}-z_{0},\hat{z}_{k}-\hat{z}_{k+1}\rangle+\langle z_{0}-z,\hat{z}_{k}-\hat{z}_{k+1}\rangle.$ (14)
We first bound the fist term of the right hand side of (2.1). By Step 2 in Algorithm 1, we get
$F(\hat{z}_{k})+DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})+6H_{k}\|z_{k+1}-\hat{z}_{k}\|(z_{k+1}-\hat{z}_{k})=0,$
$\|F(z_{k+1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\|\leqslant\frac{H_{k}}{2}\|z_{k+1}-\hat{z}_{k}\|^{2}.$
Then we obtain
$~{}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle$
$=$ $~{}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\rangle$
$~{}+\langle z_{k+1}-\hat{z}_{k+1},F(\hat{z}_{k})+DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\rangle$
$\leqslant$ $~{}\frac{H_{k}}{2}\|z_{k+1}-\hat{z}_{k+1}\|\|z_{k+1}-\hat{z}_{k}\|^{2}-6H_{k}\|z_{k+1}-\hat{z}_{k}\|\langle z_{k+1}-\hat{z}_{k+1},z_{k+1}-\hat{z}_{k}\rangle.$ (15)
@@
$~{}\lambda_{k+1}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle$
@@
where the third inequality is by the Cauchy-Schwarz inequality.
@@
$~{}\lambda_{k+1}\langle z_{k+1}-z,F(z_{k+1})\rangle$ (18)
@@
The proof is then completed by summing the above inequality for $k$ from $1$ to $s$ and using $\hat{z}_{1}=z_{0}$. ∎
The following lemma further gives bound of the sequences $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ generated by Algorithm 1.
Lemma 2.
Suppose that Assumption 1 holds. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be sequences generated by Algorithm 1. Set $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then for any $z$ and $s\geqslant 1$, we have
@@
$\|z_{s+1}-z^{*}\|\leqslant 7\|z_{0}-z^{*}\|,$ $~{}\|\hat{z}_{s+1}-z^{*}\|\leqslant 3\|z_{0}-z^{*}\|,$ (21)
@@
Proof.
By Lemma 1 and the Cauchy-Schwarz inequality, we have
@@
00 0 $\leqslant~{}\sum_{k=1}^{s}\lambda_{k+1}\langle z_{k+1}-z^{*},F(z_{k+1})\rangle+\frac{1}{24}\sum_{k=1}^{s}\|z_{k+1}-\hat{z}_{k}\|^{2}$
@@
Therefore, by the Cauchy-Schwarz inequality, we have $\|\hat{z}_{s+1}-z^{*}\|\leqslant\|\hat{z}_{s+1}-z_{0}\|+\|z_{0}-z^{*}\|\leqslant 3\|z_{0}-z^{*}\|$ and $\|z_{s+1}-z^{*}\|\leqslant\|z_{s+1}-z_{0}\|+\|z_{0}-z^{*}\|\leqslant 7\|z_{0}-z^{*}\|$. The proof is then completed. ∎

### v2

@@ 1.2 Contributions
We propose a Lipschitz-free cubic regularization (LF-CR) algorithm for solving convex-concave minimax optimization problems without knowledge of the Lipschitz constant. It can be proved that the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution with respect to the restricted primal-dual gap (resp. gradient norm) is upper bounded by $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ (resp. $\mathcal{O}(\rho\|z_{0}-z^{*}\|^{2}\epsilon^{-1})$), where $z_{0}=(x^{0},y^{0})$ is a pair of initial points, $z^{*}=(x^{*},y^{*})$ is a pair of the optimal solutions, and $\rho$ is the Lipschitz constant. This complexity bound is consistent with existing second-order methods in Lin et al. (2022). Compared to the adaptive second-order optimistic method (ASOM) in Jiang et al. (2024), which has an iteration complexity of $\mathcal{O}\left((\ell^{4/3}\rho^{2/3}\|z_{0}-z^{*}\|^{2/3}+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-2/3}\right)$ $\left(\text{resp.}\leavevmode\nobreak\ \mathcal{O}\left((\ell\rho\|z_{0}-z^{*}\|+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-1}\right)\right)$ in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap (resp. gradient norm), the LF-CR algorithm has better iteration complexity bound with respect to the Lipschitz constant $\rho$.
@@
We further propose a fully parameter-free cubic regularization (FF-CR) algorithm without any knowledge of the problem parameters, including the Lipschitz constants and the upper bound of the distance from the initial point to an optimal solution. The iteration complexity to obtain an $\epsilon$-optimal solution with respect to the gradient norm can be proved to be upper bounded by $\mathcal{O}\big{(}\rho^{2/3}\|z_{0}-z^{*}\|^{4/3}\epsilon^{-2/3}\big{)}$. Numerical experiments show the efficiency of the two proposed algorithms.
@@
To the best of our knowledge, the proposed FF-CR algorithm is a completely parameter-free second-order algorithm, and its iteration complexity is currently the best in terms of $\epsilon$ under the termination criterion of the gradient norm. Furthermore, our method only requires a mild assumption that the Hessian is Lipschitz continuous and the norm of Hessian matrix at the the initial point is upper bounded, without making additional assumptions of error bound conditions as in (Huang et al., 2022). Compared to the restarted higher-order mirror-prox algorithm with the knowledge of parameters (Ostroukhov et al., 2020), we improve the complexity for the convex-concave setting by a logarithmic factor. Moreover, our method achieves better iteration complexity compared to the adaptive second-order optimistic method (ASOM) in (Jiang et al., 2024), which has an iteration complexity of $\mathcal{O}\left((\ell\rho\|z_{0}-z^{*}\|+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-1}\right)$ in terms of gradient norm. Table 1 provides a detailed comparison of the two proposed algorithms with existing parameter-free second-order algorithms for solving convex-concave minimax problems.
Table 1: Comparison of parameter-free second-order algorithms for solving convex-concave minimax problems
Algorithms Termination Criterion 1 Iteration Complexity2
ReNewton (Liu and Luo, 2022) $\textsc{gap}(x,y;\beta)\leqslant\epsilon$ $\mathcal{O}(\rho^{2/3}D^{2}\epsilon^{-2/3})$
OSOM (Jiang and Mokhtari, 2022) $\textsc{gap}(x,y;\beta)\leqslant\epsilon$ $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$
ASOM (Jiang et al., 2024) $\textsc{gap}(x,y;\beta)\leqslant\epsilon$ $\mathcal{O}\left((\ell^{4/3}\rho^{2/3}\|z_{0}-z^{*}\|^{2/3}+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-2/3}\right)$
$\|\nabla f(x,y)\|\leqslant\epsilon$ $\mathcal{O}\left((\ell\rho\|z_{0}-z^{*}\|+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-1}\right)$
LF-CR (This paper) $\textsc{gap}(x,y;\beta)\leqslant\epsilon$ $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$
$\|\nabla f(x,y)\|\leqslant\epsilon$ $\mathcal{O}(\rho\|z_{0}-z^{*}\|^{2}\epsilon^{-1})$
FF-CR (This paper) $\|\nabla f(x,y)\|\leqslant\epsilon$ $\mathcal{O}\big{(}\rho^{2/3}\|z_{0}-z^{*}\|^{4/3}\epsilon^{-2/3}\big{)}$
  • 1
    $\textsc{gap}(x,y;\beta)$” denotes the (restricted) primal-dual gap defined in (12).
  • 2
    DD” denotes the diameter of the constraint set, $\ell$” and $\rho$” represent the Lipschitz constants of the gradient and the Hessian, respectively.
@@ 1.3 Organization
$DF(z)=\begin{bmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \nabla_{xx}^{2}f(x,y)&\leavevmode\nobreak\ \leavevmode\nobreak\ \nabla_{xy}^{2}f(x,y)\\ -\nabla_{yx}^{2}f(x,y)&-\nabla_{yy}^{2}f(x,y)\end{bmatrix}\in\mathbb{R}^{(m+n)\times(m+n)}.$ (3)
@@ 2 A Lipschitz-Free Cubic Regularization (LF-CR) Algorithm
In this section, we propose a Lipschitz-free cubic regularization (LF-CR) algorithm that does not require prior knowledge of the Lipschitz constant. We iteratively approximate the Lipschitz constants and show that the iteration complexity of the proposed algorithm remains the same order in the absence of information about the Lipschitz constants. Moreover, the LF-CR algorithm is a key subroutine in our second parameter-free algorithm that we will present in the next section.
The proposed LF-CR algorithm is based on the Newton-MinMax algorithm (Lin et al., 2022), which consists of the following two important algorithmic components at each iteration:
@@
    $\hat{z}_{k+1}=\hat{z}_{k}-\lambda_{k+1}F(z_{k+1}),$
    where $\lambda_{k+1}=\frac{c_{k}}{\rho\|z_{k+1}-\hat{z}_{k}\|}$ for some constant $c_{k}$. Note that $\|z_{k+1}-\hat{z}_{k}\|=0$ if and only if $\|F(\hat{z}_{k})\|=0,$ which will be proved later.
Although the Newton-MinMax algorithm achieves the best known iteration complexity, it requires knowing $\rho$, the Lipschitz constant of the Hessian of $f(x,y)$, which limits its applicability. To eliminate the need for the prior knowledge of $\rho$, we propose a LF-CR algorithm that employs a line search strategy in the first gradient update step. Similar line search strategies have been used in (Liu and Luo, 2022; Nesterov and Polyak, 2006; Nesterov, 2015, 2018). More detailedly, in each iteration of the proposed LF-CR algorithm, we aim to search a pair $(H_{k,i_{k}},z_{k,i_{k}})$ by backtracking such that
$\|F(z_{k,i_{k}})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k,i_{k}}-\hat{z}_{k})\|\leqslant\frac{H_{k,i_{k}}}{2}\|z_{k,i_{k}}-\hat{z}_{k}\|^{2},$ (5)
where $z_{k,i_{k}}$ is a solution of the following nonlinear equation:
$F(\hat{z}_{k})+DF(\hat{z}_{k})(z-\hat{z}_{k})+6H_{k,i_{k}}\|z-\hat{z}_{k}\|(z-\hat{z}_{k})=0.$ (6)
We then introduce how to solve the nonlinear equations in (6). Similar to the discussion in (Lin et al., 2022), (6) can be reformulated as finding a pair of $(\theta_{k,i_{k}},z_{k,i_{k}})$ such that
$\theta_{k,i_{k}}$ $=6H_{k,i_{k}}\|z_{k,i_{k}}-\hat{z}_{k}\|,$
$(DF(\hat{z}_{k})$ $+\theta_{k,i_{k}}I_{m+n})(z_{k,i_{k}}-\hat{z}_{k})=-F(\hat{z}_{k}).$
Solving the above nonlinear system of equations is equivalent to first solving the following univariate nonlinear equation with respect to $\theta$:
$\phi(\theta):=\sqrt{\|(DF(\hat{z}_{k})+\theta I_{m+n})^{-1}F(\hat{z}_{k})\|^{2}}-\frac{\theta}{6H_{k,i_{k}}}=0.$ (7)
Similar to the proof of Proposition 4.5 in (Lin et al., 2022), when $\theta>0$, $\phi(\theta)$ is a strictly decreasing and convex function. Therefore, we can first use Newton’s method to solve the one-dimensional nonlinear equation (7) to find the optimal solution of $\theta$. Then by substituting it into the following equation and solving the following linear system,
@@
we obtain $z_{k,i_{k}}$. Next, the proposed LF-CR algorithm performs an extra gradient update on $z_{k,i_{k}}$. The detailed algorithm is formally described in Algorithm 1.
@@
Step 1: Input $z_{0},\tilde{z}_{0}$; Set $H_{0}=\frac{\|DF(z_{0})-DF(\tilde{z}_{0})\|}{\|z_{0}-\tilde{z}_{0}\|}(z_{0}\neq\tilde{z}_{0})$, $k=1$, $\hat{z}_{1}=z_{1}=z_{0}.$
@@
    $H_{k}=H_{k,i}$, $z_{k+1}=z_{k,i}$; otherwise, set $H_{k,i+1}=2H_{k,i}$, $i=i+1,$ go to Step 2(b).
Step 3: If $\|z_{k+1}-\hat{z}_{k}\|=0$, stop and output $(\hat{z}_{k},H_{k})$.
Step 4: Update $\hat{z}_{k}$:
$\hat{z}_{k+1}=\hat{z}_{k}-\lambda_{k+1}F(z_{k+1}),$ (11)
    with $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$.
Step 5: If some stationary condition or stopping criterion is satisfied, let $\bar{z}_{k}=\frac{1}{\sum_{i=1}^{k}\lambda_{i+1}}\sum_{i=1}^{k}\lambda_{i+1}z_{i+1}$ and output $(\bar{z}_{k},H_{k})$; otherwise, set $k=k+1,$ go to Step 2(a).
In the following subsection, we establish the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the restricted primal-dual gap and gradient norm. Our complexity analysis builds upon the framework of (Lin et al., 2022) while addressing the more challenging scenario where the Hessian Lipschitz constant is unknown. We relax the assumption and only require that the Hessian matrix is locally Lipschitz continuous. The key difference in the proof is that we establish a lower bound for the adaptive step size $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$, where $H_{k}$ is estimated dynamically via line search rather than being a fixed constant. This requires constraining $H_{k}$ in the algorithm design and carefully considering the overhead of line search in the complexity analysis. In addition, under the assumption of local Lipschitz continuity, we need to prove the boundedness of the sequence of the iteration points.
@@ 2.1 Complexity analysis
[…]
Definition 1
A pair $(x^{*},y^{*})$ is called an optimal solution of problem (1), if
$f(x^{*},y)\leqslant f(x^{*},y^{*})\leqslant f(x,y^{*})\leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ x\in\mathbb{R}^{m}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ y\in\mathbb{R}^{n}.$
Similar to (Lin et al., 2022), we define the restricted primal-dual gap function for (1) as follows.
Definition 2
@@
[…]
Definition 3
[…]
Definition 4
[…]
Similar to the definition of gradient Lipschitz continuity in (Vladarean et al., 2021), we introduce the following definitions for local and global Hessian Lipschitz continuity.
Definition 5
A function $f(x,y)$ is globally Hessian Lipschitz continuous, if there exists a positive scalar $\rho$ such that
$\|\nabla^{2}f(x,y)-\nabla^{2}f(x^{\prime},y^{\prime})\|\leqslant\rho\|(x-x^{\prime},y-y^{\prime})\|,\leavevmode\nobreak\ \forall(x,y),(x^{\prime},y^{\prime})\in\mathbb{R}^{m}\times\mathbb{R}^{n}.$
Definition 6
A function $f(x,y)$ is locally Hessian Lipschitz continuous, if there exists a positive scalar $\rho({\mathcal{C}})$ such that for any compact subset $\mathcal{C}\in\mathbb{R}^{m}\times\mathbb{R}^{n}$,
$\|\nabla^{2}f(x,y)-\nabla^{2}f(x^{\prime},y^{\prime})\|\leqslant\rho({\mathcal{C}})\|(x-x^{\prime},y-y^{\prime})\|,\leavevmode\nobreak\ \forall(x,y),(x^{\prime},y^{\prime})\in\mathcal{C}.$
It is important to note that the local Hessian-Lipschitz condition is strictly weaker than the global condition, and therefore covers a wider range of functions. For example, functions like $e^{x}$ and $x^{p}\leavevmode\nobreak\ (p\geqslant 4)$ are locally Hessian-Lipschitz continuous, but not globally.
We also need to make the following assumptions about the function $f(x,y).$
Assumption 1
The function $f(x,y)$ is convex-concave, i.e., $f(\cdot,y)$ is convex for any fixed $y\in\mathbb{R}^{n}$ and $f(x,\cdot)$ is concave for any fixed $x\in\mathbb{R}^{m}$. The optimal solution of problem (1) exists.
Assumption 2
The function $f(x,y)$ is locally Hessian Lipschitz continuous.
Lemma 1
Suppose that Assumptions 1 and 2 hold. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be the sequences generated by Algorithm 1. Then, $\|z_{k+1}-\hat{z}_{k}\|=0$ if and only if $\|F(\hat{z}_{k})\|=0.$
Proof  By Step 2 in Algorithm 1, we get
$F(\hat{z}_{k})+DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})+6H_{k}\|z_{k+1}-\hat{z}_{k}\|(z_{k+1}-\hat{z}_{k})=0,$ (13)
$\|F(z_{k+1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\|\leqslant\frac{H_{k}}{2}\|z_{k+1}-\hat{z}_{k}\|^{2}.$ (14)
Then, on the one hand, if $\|z_{k+1}-\hat{z}_{k}\|=0$, then it is straightforward to prove $\|F(\hat{z}_{k})\|=0$. On the other hand, if $\|F(\hat{z}_{k})\|=0$, by taking the inner product of both sides of (13) with $z_{k+1}-\hat{z}_{k}$, we obtain
$\left\langle z_{k+1}-\hat{z}_{k},DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\right\rangle+6H_{k}\|z_{k+1}-\hat{z}_{k}\|^{3}=0,$
which implies $\|z_{k+1}-\hat{z}_{k}\|=0$. The proof is completed.
@@
Lemma 2
Suppose that Assumptions 1 and 2 hold. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be the sequences generated by Algorithm 1. Assuming that $\|z_{k+1}-\hat{z}_{k}\|\neq 0$, set $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If in addition $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then for any $z\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ and $s\geqslant 1$, we have
@@
Proof  By (11) in Algorithm 1, we have
$\leavevmode\nobreak\ \lambda_{k+1}\langle z_{k+1}-z,F(z_{k+1})\rangle$
$=$ $\leavevmode\nobreak\ \lambda_{k+1}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle+\lambda_{k+1}\langle\hat{z}_{k+1}-z_{0},F(z_{k+1})\rangle+\lambda_{k+1}\langle z_{0}-z,F(z_{k+1})\rangle$
$=$ $\leavevmode\nobreak\ \lambda_{k+1}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle+\langle\hat{z}_{k+1}-z_{0},\hat{z}_{k}-\hat{z}_{k+1}\rangle+\langle z_{0}-z,\hat{z}_{k}-\hat{z}_{k+1}\rangle.$ (16)
We first bound the first term of the right-hand side of (2.1). By (13) and (14), we obtain
$\leavevmode\nobreak\ \langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle$
$=$ $\leavevmode\nobreak\ \langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\rangle$
$\leavevmode\nobreak\ +\langle z_{k+1}-\hat{z}_{k+1},F(\hat{z}_{k})+DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\rangle$
$\leqslant$ $\leavevmode\nobreak\ \frac{H_{k}}{2}\|z_{k+1}-\hat{z}_{k+1}\|\|z_{k+1}-\hat{z}_{k}\|^{2}-6H_{k}\|z_{k+1}-\hat{z}_{k}\|\langle z_{k+1}-\hat{z}_{k+1},z_{k+1}-\hat{z}_{k}\rangle.$ (17)
@@
$\leavevmode\nobreak\ \lambda_{k+1}\langle z_{k+1}-\hat{z}_{k+1},F(z_{k+1})\rangle$
@@
where the last inequality is by the Cauchy-Schwarz inequality.
@@
$\leavevmode\nobreak\ \lambda_{k+1}\langle z_{k+1}-z,F(z_{k+1})\rangle$ (20)
@@
The proof is then completed by summing the above inequality for $k$ from $1$ to $s$ and using $\hat{z}_{1}=z_{0}$.
The following lemma further gives bounds for the sequences $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ generated by Algorithm 1.
Lemma 3
Suppose that Assumptions 1 and 2 hold. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be the sequences generated by Algorithm 1. Assuming that $\|z_{k+1}-\hat{z}_{k}\|\neq 0$, set $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If in addition $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then for any $z\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ and $s\geqslant 1$, we have
@@
$\|z_{s+1}-z^{*}\|\leqslant 7\|z_{0}-z^{*}\|,$ $\leavevmode\nobreak\ \|\hat{z}_{s+1}-z^{*}\|\leqslant 3\|z_{0}-z^{*}\|,$ (23)
@@
Proof  By Lemma 2 and the Cauchy-Schwarz inequality, we have
@@
$0$ $\leqslant\leavevmode\nobreak\ \sum_{k=1}^{s}\lambda_{k+1}\langle z_{k+1}-z^{*},F(z_{k+1})\rangle+\frac{1}{24}\sum_{k=1}^{s}\|z_{k+1}-\hat{z}_{k}\|^{2}$
@@
Therefore, by the Cauchy-Schwarz inequality, we have $\|\hat{z}_{s+1}-z^{*}\|\leqslant\|\hat{z}_{s+1}-z_{0}\|+\|z_{0}-z^{*}\|\leqslant 3\|z_{0}-z^{*}\|$ and $\|z_{s+1}-z^{*}\|\leqslant\|z_{s+1}-z_{0}\|+\|z_{0}-z^{*}\|\leqslant 7\|z_{0}-z^{*}\|$. The proof is then completed.